June 25, 2020 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Time: 11:00 a.m. JLab time

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

  1. Checks of monitoring comments of Batch 5-8
    1. Timing anomalies: 72123-72416 (2ns peaks), 72417-72435 (time offset) -> Mark
    2. Hit efficiencies: 72316-72435 (especially visible in "enhanced" efficiencies) -> related to next item?
    3. Recon Matching: 72271 (inefficiency at module~12, and 90deg), 72385 (degraded efficiency at module~47) -> Elton
  2. Long term items

Tentative Agenda

  1. Announcements:
  2. Action Items
  3. Run Updates
    1. FCAL
    2. BCAL
  4. Calibrations
    1. BCAL Gain Calibrations for different Batches (B2-B8) Update
      1. BCAL gain stability across all channells for the Batches (2-8) ChannelByChannelMass
  5. Monitoring
    1. BCAL LED [2]
    2. BCAL (offline)
    3. FCAL (offline) status
  6. Reconstruction
  7. Efficiencies
    1. Photon Reconstruction Systematics Task Force
  8. Simulations
  9. Any other business


Attending: Matt, Beni. Sasha, Christina, Churamani, Colin, Edmundo, Curtis, Eugene, Jon, Karthik, Rebecca, Mark, Sean, Varun, Susan, Zisis

  1. Announcements: Karthik and Varun passed the written component of the PhD Comprehensive exam.
  2. Action Items
    1. Mark looked at item 1. Only small offset changes. Towards the end of the run (~ 2-day subset of batch 8) the trigger changed and so did the efficiency. CCAL was readout but maybe not in the trigger -- to be checked.
  3. Run Updates
    1. FCAL (Colin): FCAL has been on for ~1 week; thanks to Mark and Nick. Bases are being monitored; 4 bases gave HV comm alarms. A 'preemptive strike' for base repairs was discussed at length. In the end, Colin recommended a conservative approach in not swappnig out bases before the run and keeping ~10 bases on hand for replacement during the run. After the run, a comprehensive look and action for bases (HV, comm, high current draw, coax cable) both within the 3 inner rings, but also elsewhere in the FCAL can be pursued. Fernando has stated in the past that he knows of no vendor that can make new boards.
    2. BCAL: it is on and operatting ok; currently the SiPMs are warm (ambient temperature) and will be cooled ~1 week before the run starts. Varun and Jon will take LED measurements before and after the run, to add to our past results on radiation damage.
  4. Calibrations
    1. BCAL Gain Calibrations for different Batches (Karthik): Karthik showed th pi0 Mass (after nonlinear energy corrections) as a function of Run Number. Overall he sees a 0.4% drift downwards as a function of run number, similar to what Varun sees with the LEDs. Some junk runs leaked into the skims and those will be removed. A cluster of runs in batch 7 jumped up, and this will be investigated. Note that batch 2 gains are used for batches 2-8. The gains per batch will be examined. The non-linearity correction plot was updated using the new symmetric condition: Eg1-Eg2<10%. New plots showing pi0 mass stability as a function of layer were shown. Mod 26 (after it got hit) sticks out; its neighbours also show an effect (Mods 25, 27, 28). Batch 4 was change by Sean so that the incident is at the batch boundary. Mark recommends calibrating before and after the incident. Beni suggested to combine batches by trends seen in Karthik's slide 1. Mark: these plots hide individual channel behaviour, and therefore radiation damage effects. Jon states that any damage is still a long way from affecting the trigger and resolution.
    2. BCAL gain stability across all channells for the Batches (Karthik): Karthik's send set of slides showed the gain stability per channel.
  5. Monitoring
    1. BCAL LED (Varun): Varun looked the the LED monitoring for batches 4-8. No major anomalies. In batch 6 the LED pulser had an issue, and Mark and Elton rebooted the VME crate which fixed it. Batch 7 has gaps because some of those runs were test runs.
    2. BCAL (offline): Tolga sent regrets before the meeting; he was following Elton's talk on scintillators.
    3. FCAL (offline): Susan reviewed the monitoring histos for batches 5-8. She summarized the observations in a table. Batch 5 looked near perfect and therefore had no comments. Observations: FCAL Local Hit Time RMS: wider for outer 'rings' ('radius' > 15) and structure upper left quadrant egde (batch 6). Also, FCAL Neutral Shower dt vs Energy: vertical bands appear; some are likely due to LEDs in the BCAL and should be filtered from the skims by imposing a criterion like Foda did: more than 1200 channels firing means LED; hard to see how LEDs whould do the same in the FCAL though. Matt suggested to look at the possibility of pedestal drift. Slide 7: Batch 8 showed a dramatic extra RF peak in FCAL Neutral Shower dt. Sean: CCAL trigger should be removed from monitoring (runs 72416-72428). Slide 8: dt(FCAL-RF) shifted by -8ns. Slide 10: FCAL Clusters Phi different (high intensity scan: compare to Karthik?). Matt comment that the high rate LED data causes a shift in the pi0 mass.
  6. Reconstruction: no report.
  7. Efficiencies:
    1. Photon Reconstruction Systematics Task Force: Matt reported that Rebecca will continue Jon's efficiency study using omega skims. Will validate the shower quality in data and MC for omega events. All past efficiency studies will be redone, after all changes in reconstruction and MC.
  8. Simulations: no report.
  9. Any other business
    1. Beni discussed the splitoffs and cluster merging as part of the multiplicity task force and will present something at the next CALWG meeting. Such effects depend on event topology. Matt: at high energy for pi0 the two photons anyhow come close to each other; hadronic splitoffs are messier in FCAL, less so in BCAL. Mark: Wil MCGinley's clusterizer is pretty sophisticated. Jon: 5% of the time 7 showers appear in eta->3pi0 channel. Sean: use the event viewer to look at such issues. Susan: study the opening angle like at WASA (Christine also suggested this in the past). Sean: Define the problem in (theta,phi) plane. Matt: well below the mass of the pi0 another peak appears sometimes. Eugene: quantify good fiducials cuts. Note that Dave M and Jane had studied splitoffs a while back. We should retrieve their work.